For a set of data, the coefficient of correlation is r=0.57 and the variable X and Y are directly related. The coefficient of determination is r equal to
1. 0,75
2. 0,57
3. 0,32
4. 0,49
5. 0,67
1
Expert's answer
2020-11-11T19:30:00-0500

Solution:

Coeficient of determination (R squared) can be calculated as:

"R^2=r^2=(0.57)^2=0.3249"

Answer:

3. 0,32